OpenStudy (anonymous):

(cscy+coty)(cosy-coty)/(cscy) How does this simplify? I tried to expand it and work it out and got that this simplified is cscy. Am I correct?

OpenStudy (helder_edwin):

\[\large \frac{(\csc y+\cot y)(\cos y-\cot y)}{\csc y}= \] \[\large =\frac{(\frac{1}{\sin y}+\frac{\cos y}{\sin y})(\cos y-\frac{\cos y}{\sin y})} {\frac{1}{\sin y}} \]

OpenStudy (anonymous):

Woah I was way off...I'm really confused with these identities

OpenStudy (helder_edwin):

\[\large \frac{\frac{1}{\sin y}(1+\cos y)\cos y (1-\frac{1}{\sin y})}{\frac{1}{\sin y}} \] \[\large =(1+\cos y)\frac{\cos y}{\sin y}(\sin y-1) \]

OpenStudy (helder_edwin):

i think this is right so far. sorry i gotta go.

OpenStudy (anonymous):

its okay thank you!

OpenStudy (loser66):

@megprang show me your work, please. I don't know how to get csc y like yours. hehe

OpenStudy (anonymous):

Based on what he did up there I think I completely got it wrong. I pretty much just tried (hyp./opp.)+(adj./hyp.) times (hyp./opp.)-(adj./opp.) and then divided that by hyp./opp. I assumed that the top would cancel eachother out and that I would be left with the denominator which is the hyp./opp. which is equal to the cosecant.